# Demographic-aware fine-grained classification of pediatric wrist fractures

### v3

@@ Abstract
Wrist pathologies are frequently observed, particularly among children who constitute the majority of fracture cases. Computer vision presents a promising avenue, contingent upon the availability of extensive datasets, a notable challenge in medical imaging. Therefore, reliance solely on one modality, such as images, proves inadequate, especially in an era of diverse and plentiful data types. In this study, we employ a multifaceted approach to address the challenge of recognizing wrist pathologies using an extremely limited dataset. Initially, we approach the problem as a fine-grained recognition task. Secondly, we enhance network performance by fusing patient metadata with X-rays. Thirdly, we improve the performance further by utilizing weights trained on a separate fine-grained dataset. While metadata integration has been used in other medical domains, this is a novel application for wrist pathologies.
@@ 1 Introduction
Despite this progress, metadata integration has not been investigated for wrist pathology classification. This omission is critical because wrist anatomy and pathology presentation are strongly age- and sex-dependent. Carpal ossification occurs in well-defined stages, the distal radial and ulnar physes close at different ages for boys and girls [9, 10], and common pediatric fracture types (e.g., buckle, greenstick) differ from adult patterns [11, 12]. Without such demographic context, radiographic findings may be misinterpreted as abnormal when they are not, or vice versa.
To address this gap, we extend metadata integration to musculoskeletal imaging, framing wrist abnormality recognition as a fine-grained visual recognition (FGVR) task. Unlike prior multimodal efforts that primarily used conventional CNNs, we fuse demographic metadata with radiographs to capture subtle, age- and sex-specific morphological cues that image-only models often overlook. Building on our prior work image-only fine-grained recognition work [13, 14], we propose a joint framework for pediatric wrist pathology analysis and hypothesize that it will outperform image-only and CNN-based baselines.
@@ 2.1 Dataset Curation & Patient Metadata
We curated two datasets from GRAZPEDWRI-DX [15]. (1) A “limited” set of three wrist pathologies (fracture, bone anomaly, soft tissue), excluding the “metal” class, following our prior work [13, 14]. (2) A binary “Fracture vs. No Fracture” dataset using all images, with the non-fracture class augmented for balance. Final splits are shown in Table 1. The cohort included 6,091 patients (2,688 female, 3,402 male, ages 0.2–19 years, mean 10.9). Patient metadata is derived from the filename structure of each image. Fig. 2 shows distributions: fractures are more common in males, while anomalies concentrate in older groups.
@@ 2.2 Architectural Details
This work utilizes a hybrid architecture that integrates both visual data and meta-information, drawing inspiration from the architecture called “MetaFormer” [16].
Visual Backbone. As depicted in Fig. 1, the architecture is organized into five stages. The first stage (S0) begins with a 3-layer convolutional stem that processes raw wrist images, introducing locality and inductive bias for edges and textures. Stages S1 and S2 consist of MBConv blocks enhanced with squeeze-and-excitation, which increase representational power by modeling inter-channel dependencies. These stages progressively downsample the feature maps while expanding channel dimensionality, ensuring the network can capture both low-level bone structures and mid-level soft tissue patterns. The final two stages, S3 and S4, transition from convolution to transformer-based blocks with relative positional encoding. This design allows the model to capture global dependencies, which are crucial for subtle pathological differences that span larger contexts (e.g., comparing radius vs. ulna contours or detecting diffuse soft tissue swelling).
Downsampling at the beginning of each stage reduces resolution while maintaining contextual continuity. In S3 and S4, overlapping patch embedding is employed via strided convolution with zero-padding [17], which provides partially redundant tokens that preserve boundary details, important when localizing hairline fractures or subtle cortical irregularities. Table 2 outlines the configuration details of the three variants of the architecture used in our experiments.
@@
Transformer Fusion. Standard self-attention lacks spatial awareness; therefore, we incorporate a relative position bias. Specifically, a bias matrix $B\in\mathbb{R}^{((M^{2}+N)\times(M^{2}+N))}$ is added during attention computation:
@@
Here, $Q$, $K$, and $V$ are query, key, and value matrices of size $\mathbb{R}^{((M^{2}+N)\times d)}$. $M^{2}$ corresponds to the number of image patches and $N$ is the count of additional tokens (class and meta). This positional encoding ensures that the model attends appropriately to the relative arrangement of wrist bones, where spatial context is essential for distinguishing true pathology from normal variants.
The Relative Transformer Block is composed of multi-head self-attention (MSA), layer normalization (LN), and a multi-layer perceptron (MLP). It operates on a token sequence $z_{0}$ that combines visual, meta, and class tokens:
@@
$z^{\prime}_{i}$ $=\text{MSA}(\text{LN}(z_{i-1}))+z_{i-1}$
@@
This formulation enables the network to learn correlations not only within visual patches but also across modalities. For example, the model can attend jointly to a distal radius patch and the token representing a 70-year-old female, thereby weighting its decision toward fracture likelihood based on both visual evidence and prior epidemiology.
Multi-Scale Aggregation. Stages S3 and S4 output separate class tokens, $z^{1}_{\text{class}}$ and $z^{2}_{\text{class}}$, which encode fused visual and meta representations at different scales. Since dimensions differ, $z^{1}_{\text{class}}$ is projected using an MLP to align with $z^{2}_{\text{class}}$. These are then concatenated and passed through a 1D convolution followed by normalization:
@@
The final output $y$ aggregates both global image evidence and auxiliary patient priors, which is especially important in fine-grained medical classification.
Meta-information Integration. When external attributes are available, they are embedded into the same feature space as vision tokens. In our case, age and gender are encoded as low-dimensional vectors and passed through a non-linear fully connected embedding function $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{d}$. This ensures that demographic priors, e.g., higher fracture prevalence in elderly females, are directly fused with image tokens. Importantly, meta-information is gradually incorporated during training to prevent early over-reliance on attributes, allowing the visual encoder to remain robust.
@@ 3.1 Baseline Comparison
Table 3 shows that MetaFormer (79.4%) outperforms widely used CNNs such as ResNet50 (77.7%) and DenseNet201 (76.8%). This suggests that transformer-based architectures with fine-grained design are better suited to capture subtle morphological cues in pediatric wrist radiographs, where differences between normal anatomy and pathology can be minimal, giving credence to our previous works [13, 14].
@@ 3.2 Metadata Integration & Pretraining
Integrating demographic metadata (age, sex) consistently improved performance (Table 4). Both early and late fusion surpassed vision-only models, with late fusion achieving the highest gains in two out of three configurations. This finding aligns with the intuition that age and sex provide essential developmental context, particularly relevant in pediatrics, where normal ossification patterns vary across individuals. Attribute-level analysis (Table 6) further indicates that both age and sex independently contribute to performance, but their combination yields the largest benefit, supporting the clinical relevance of demographic context.
The best configuration “FG-2-inat” with fine-grained pretraining reached 81.4%, a 1.5% improvement over image-only training. Fine-grained pretraining on iNaturalist [27] improved generalization beyond ImageNet pretraining, reflecting the value of transferring from domains where subtle inter-class variations dominate.
@@
Fig. 3: Heatmaps generated from the final convolutional layer in stage 2 (S2) of our configuration fine-grained-FG-2 on a limited dataset.

### v4

@@ Abstract
Wrist pathologies are frequently observed, particularly among children who constitute the majority of fracture cases. Computer vision presents a promising avenue, contingent upon the availability of extensive datasets, a notable challenge in medical imaging. Therefore, reliance solely on one modality, such as images, proves inadequate, especially in an era of diverse and plentiful data types. This study addresses the problem using a multifaceted approach: framing it as a fine-grained recognition task, fusing patient metadata with X-rays, and leveraging weights from a separate fine-grained dataset rather than from a coarse-grained dataset like ImageNet. Unlike prior work, this is the first application of metadata integration for wrist pathology recognition. Our results show that combining fine-grained transformer approach, fine-grained pre-training, and metadata integration improves diagnostic accuracy by 2% on small custom curated dataset and over 10% on a larger fracture dataset.
@@ 1 Introduction
Despite this progress, metadata integration has not been investigated for wrist pathology classification. This omission is critical because wrist anatomy and pathology presentation are strongly age- and sex-dependent. Carpal ossification occurs in well-defined stages, the distal radial and ulnar physes close at different ages for boys and girls [9, 10], and common pediatric fracture types (e.g., buckle, greenstick) differ from adult patterns [11, 12]. Without age/sex context, a model might misclassify normal developmental features (like unfused growth plates) as abnormalities.
To address this gap, we extend metadata integration to musculoskeletal imaging. We first frame wrist abnormality recognition as a fine-grained visual recognition (FGVR) task, unlike prior multimodal efforts that primarily used conventional CNNs. On top of this, we fuse demographic metadata with radiographs to capture subtle, age- and sex-specific morphological cues that image-only models often overlook. Building on our prior work image-only fine-grained recognition work [13, 14], we propose a joint framework for pediatric wrist pathology analysis and hypothesize that it will outperform image-only and CNN-based baselines.
@@ 2.1 Dataset Curation & Patient Metadata
We curated two datasets from GRAZPEDWRI-DX [15]. (1) A “limited” set of three wrist pathologies (fracture, bone anomaly, soft tissue), excluding the “metal” class, following our prior work [13, 14]. (2) A binary “Fracture vs. No Fracture” dataset using all images, with the non-fracture class augmented for balance. Final splits are shown in Table 1. The cohort included 6,091 patients (2,688 female, 3,402 male, ages 0.2–19 years, mean 10.9). Fig. 1 shows distributions: fractures are more common in males, while anomalies concentrate in older groups.
@@ 2.2 Architectural Details
This work utilizes a hybrid architecture that integrates both visual data and meta-information, drawing inspiration from the architecture called “MetaFormer” [16]. As depicted in Fig. 2, the architecture is organized into five stages. The first stage (S0) starts with a basic 3-layer convolutional stem, which processes raw input images. Subsequent stages progressively downsample the feature maps and increase channel dimensionality. Stages S1 and S2 consist of MBConv blocks enhanced with squeeze-and-excitation mechanism. The final two stages, S3 and S4, transition to transformer-based blocks that employ relative positional encoding. This design allows the model to capture global dependencies, which are crucial for subtle pathological differences that span larger contexts (e.g., comparing radius vs. ulna contours or detecting diffuse soft tissue swelling).
Downsampling is applied at the beginning of each stage. In S3 and S4, overlapping patch embedding is implemented using convolutions with a stride of 2 and zero-padding, effectively reducing resolution while maintaining contextual overlap [17]. Table 2 outlines the configuration details of the three variants of the architecture used in our experiments.
@@
To overcome the lack of spatial awareness in standard self-attention, the model incorporates a relative position bias. Specifically, a bias matrix $B\in\mathbb{R}^{((M^{2}+N)\times(M^{2}+N))}$ is added during attention computation:
@@
Here, $Q$, $K$, and $V$ are the query, key, and value matrices of size $\mathbb{R}^{((M^{2}+N)\times d)}$, where $M^{2}$ corresponds to the number of image patches and $N$ is the count of extra tokens (e.g., class and meta tokens). The bias matrix $\hat{B}$ is parameterized as described in [18], supporting relative positions from M1-M-1 to $M+1$. A shared bias is applied for non-visual tokens since their spatial position is undefined.
The Relative Transformer Block comprises multi-head self-attention (MSA), layer normalization (LN), and a multi-layer perceptron (MLP), applied to the input token sequence $z_{0}$:
@@
$z^{\prime}_{i}$ $=\text{MSA}(\text{LN}(z_{i-1}))+z_{i}$
@@
In this formulation, $z_{i}\in\mathbb{R}^{(M^{2}+N)\times d}$ captures the updated token sequence at each transformer layer.
Stages S3 and S4 each generate a class token, $z^{1}_{\text{class}}$ and $z^{2}_{\text{class}}$, which encode combined visual and meta representations. Since these tokens may differ in dimension, $z^{1}_{\text{class}}$ is first projected to match the shape of $z^{2}_{\text{class}}$ using an MLP. The two are then concatenated and passed through a 1D convolution followed by normalization:
@@
The final output $y$ aggregates information across multiple scales and token types. When using external attributes as metadata, these are encoded as vectors and embedded using a fully connected non-linear function $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{d}$. In our setup, the demographic attributes consist of age and gender. The age attribute was normalized using min–max scaling, while the gender attribute was one-hot encoded. This results in a 2D input vector, which is then projected into the same embedding space as the vision tokens, ensuring consistency between visual and meta representations.
@@ 3.1 Baseline Comparison
Table 3 shows that MetaFormer (79.4%) outperforms widely used CNNs such as ResNet50 (77.7%) and DenseNet201 (76.8%). This gives credence to the fact that transformer-based architectures with fine-grained design are better suited to capture subtle morphological cues in pediatric wrist radiographs, where differences between normal anatomy and pathology can be minimal.
@@ 3.2 Metadata Integration & Pretraining
Integrating demographic metadata (age, sex) consistently improved performance (Table 4). Both early and late fusion surpassed vision-only models, with late fusion achieving the highest gains in two out of three configurations. This finding aligns with the intuition that age and sex provide essential developmental context, particularly relevant in pediatrics, where normal ossification patterns vary across individuals. Late fusion may be more robust to noise in individual modalities, as combining information at a later stage allows the final decision to benefit from multiple sources, potentially mitigating the impact of noise or errors in any single modality.
Attribute-level analysis (Table 6) further indicates that both age and sex independently contribute to performance, but their combination yields the largest benefit, supporting the clinical relevance of demographic context.
We then selected the “FG-2” configuration and used the weights obtained from training it on the iNaturalist dataset [28] to create our “FG-2-inat” configuration. This surpassed all other configurations, whether using image data alone or in conjunction with metadata. This configuration reached 81.4%, a 1.5% improvement over image-only training. It seems that the fine-grained pretraining improved generalization beyond ImageNet pretraining, reflecting the value of transferring from domains where subtle inter-class variations dominate. Table 5 shows per class sensitivity, specificity, and precision scores for configuration “FG-2-inat”.
@@
Fig. 3: Heatmaps generated from the final convolutional layer in stage 2 (S2) of our configuration “FG-2-inat” on a limited dataset.